OpenStudy (anonymous):

Write the equation of a circle with a center at (0, 0) and a radius of 5.

Nnesha (nnesha):

(x-h) +(y - k) = r^2 equation of circle where center is (h,k) and r is radius so replace h ,k and r by their values

OpenStudy (anonymous):

The equation is (x-(x coordinate)) + (y-(y coordinate) = r^2

OpenStudy (jdoe0001):

\(\bf (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{purple}{ r}}^2 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad radius={\color{purple}{ r}}\) plug in the given values

OpenStudy (usukidoll):

ok using jdoe's formula..... and your given information (0,0) r = 5

OpenStudy (usukidoll):

now let's start on the radius... that's easy... since r = 5. r^2 =?

OpenStudy (anonymous):

25

OpenStudy (usukidoll):

yes!

OpenStudy (usukidoll):

ok... so the center is (x,y) which is (0,0) which is x = 0 and y = 0

OpenStudy (usukidoll):

\[\bf (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{purple}{ r}}^2 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad radius={\color{purple}{ r}}\] sorry sorry I used the wrong letters it's center (h,k) h = 0 k = 0 x*(

OpenStudy (usukidoll):

from (0,0) we know that h = 0 and k = 0 so we can just plug those guys into the formula

OpenStudy (anonymous):

\[(0-h)^2+(0-k)^2=25\]

OpenStudy (usukidoll):

no I made a mistake earlier... from (0,0) we know that h = 0 and k = 0 so we can just plug those guys into the formula

OpenStudy (usukidoll):

x and y should've been left alone... been dealing with that formula so much lol

OpenStudy (anonymous):

\[(x-0)^2+(y-0)^2=25\]

OpenStudy (usukidoll):

yay. we're almost done. we just have to clean it up a bit x-0 = x and y-0 = y

OpenStudy (anonymous):

so for this one.... Write the equation of a circle with a center at (1, 2) and a diameter of 15. it would be....... (x-1)^2 +(y-2)^2=15 ???

Nnesha (nnesha):

radius is half of diameter ..... r= d/2 r=radius diameter =d

OpenStudy (usukidoll):

almost correct... the left is fine. you might have to solve for the radius on this one

OpenStudy (anonymous):

so 7.5 ???

OpenStudy (anonymous):

is my radius

OpenStudy (usukidoll):

yeah, but I would leave it as 15/2. teachers these days cringe on decimals.

Nnesha (nnesha):

lol true story

OpenStudy (anonymous):

haha thanks you two :)